Ruth invest $10,000 dollarsin two accounts. one account has an annual interest rate of 7%, and the other account has an annual interest rate of 5%. Let I represent the total interest earned in one year. Then Ruth's investemnt can be modeled by the system of equations x+y=10,000 and 0.07x=0.05y=I.
a. determind a solution fo this system of equations that represents a possible investment, and find the value of I that corresponds to your solution.

Answer by stanbon(48551)

Ruth invest $10,000 dollarsin two accounts. one account has an annual interest rate of 7%, and the other account has an annual interest rate of 5%. Let I represent the total interest earned in one year. Then Ruth's investemnt can be modeled by the system of equations x+y=10,000 and 0.07x+0.05y=I.
-----------------------------------
a. determin a solution fo this system of equations that represents a possible investment, and find the value of I that corresponds to your solution.
-------------
Let amount invested at 7% be "x" ; Interest is 0.07x dollars
Let amount invested at 5% by "y" ; Interest is 0.05y dollars
--------
equations x+y=10,000 and 0.07x+0.05y=I
x=10000-y
0.07(1000-y)+0.05y =I
7000-0.02y=I
y=(7000-I)/0.02
--------------
Substitute any number you want for "I" and you can solve for "y" and then "x".
Cheers,
Stan H.